40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 5 20:37:34 2011, jamie, Summary, General, Summary of today's in-vacuum work 
    Reply  Sat Aug 6 03:10:05 2011, Suresh, Summary, General, Summary of today's in-vacuum work ETMX_OSEMs.pngAfter_osem_adjustment.JPGBefore_osem_adj.JPG
       Reply  Sat Aug 6 13:38:02 2011, rana, Summary, General, Summary of today's in-vacuum work 
          Reply  Sun Aug 7 11:10:54 2011, Suresh, Summary, General, OSEM positioning 
    Reply  Sun Aug 7 14:01:58 2011, kiwamu, Update, Green Locking, X green beam re-aligned 
       Reply  Sun Aug 7 18:41:12 2011, Jenne, Update, Green Locking, X green beam re-aligned 
Message ID: 5133     Entry time: Sun Aug 7 14:01:58 2011     In reply to: 5127     Reply to this: 5135
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: X green beam re-aligned 

[Jenne / Kiwamu]

 The X green beam has been realigned to compensate the effect of the ETMX repositioning.

After the alignment we became able to lock the 00 mode with the X green beam.

 

For the alignment:

  spot position on the ETMX mirror = within ~ 1 cm. This number is strictly constrained by a homemade aluminum iris that Jamie put last Friday.

  spot position on the ITMX mirror = unknown, but looks pretty good on the CCD camera.

  spot position on the PSL table = ~ 1 mm downward from what it used to be. The horizontal alignment is perfect.

Conclusions :

  The X green beam again became a reference of the beam axis.

  The ETMX suspension tower is in a good place.

Quote from #5127

Kiwamu will work on the green alignment over the weekend.  Assuming everything works out, we'll try the same procedure on ETMY on Monday.

 

ELOG V3.1.3-